# Notes on Bi-parameter Paraproducts

### v1

@@ 1. Introduction
This work is a sequel to our previous two papers on the boundedness properties of paraproducts [20, 21]. Here, we are concerned with sharpness of the existing bounds for the remaining dyadic bi-parameter paraproducts not studied in [20]. For various applications and properties of paraproducts, we refer the reader to [3, 2, 15, 16, 17, 14].
@@
$\pi^{\varepsilon}(f,g)(x):=\sum_{I\in\mathcal{D}}\left\langle f,h_{I}^{\varepsilon_{1}}\right\rangle\left\langle g,h_{I}^{\varepsilon_{2}}\right\rangle h_{I}^{\varepsilon_{1}+\varepsilon_{2}}(x),\quad\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\{0,1\}^{2}\backslash\{(0,0)\},\quad x\in\mathbb{R},$
where $I$ stands for a dyadic interval on the line, $\mathcal{D}$ denotes the collection of all such intervals, $h_{I}$ is the Haar wavelet associated with $I$, and $h_{I}^{\epsilon}$ is given by
$h_{I}^{\epsilon}(x):=\begin{cases}h_{I}(x),&\epsilon=1,\\ \bar{\chi}_{I}(x),&\epsilon=0,\end{cases}\quad x\in\mathbb{R},$
where $\bar{\chi}_{I}:=\frac{\chi_{I}}{|I|}$, and the sum $\varepsilon_{1}+\varepsilon_{2}$ is understood modulo 2. It is then easy to see that, for sufficiently nice functions, we have
@@
Fairly well-known arguments show that, for $0<p<\infty$, the bilinear form $\pi^{(0,1)}$ is bounded from $H^{p}_{d}(\mathbb{R})\times BMO_{d}(\mathbb{R})$ and $L^{\infty}(\mathbb{R})\times H^{p}_{d}(\mathbb{R})$ to $H^{p}_{d}(\mathbb{R})$. Here, for $0<p<\infty$, $H^{p}_{d}(\mathbb{R})$ denotes the dyadic Hardy space on the line, (quasi-)normed by
@@
It is enough to recall the square function characterization of dyadic Hardy spaces, which states that
@@
As expected, the $H_{d}^{p}(\mathbb{R})$-to-$H_{d}^{q}(\mathbb{R})$ norm of $\pi^{\prime\prime}_{g}$ is also comparable to $\|g\|_{H_{d}^{r}(\mathbb{R})}$ in the full range of exponents, when $\frac{1}{q}=\frac{1}{p}+\frac{1}{r}$. In the reflexive range, this result was established in [18] (Theorem 12.2, p. 128), and in Theorem 2.1 we give a new proof in the product setting, to which we now turn.
@@ 2. Unmixed Paraproducts
where the supremum is taken over all open sets of finite positive measure. Since it will be clear from the context, we use the same notation $M$ for the dyadic maximal operator on the line and the strong operator defined above. We also refer the reader to [20] for the reason behind the above equivalences, and to [22, 6, 9, 7] for more on Hardy spaces in the product setting. See also [5, 19]. The adjoint of this operator, which is unmixed too, is given by
@@
$\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\|S(f)\|_{L^{p}(\mathbb{R}^{2})},\quad 0<p<\infty,\quad S(f):=\Big(\sum_{R}|f_{R}|^{2}\frac{\chi_{R}}{|R|}\Big)^{\frac{1}{2}}.$
Now, the above bounds are simple consequences of the crucial pointwise inequality
@@
the square function characterization of product Hardy spaces, and Hölder’s inequality.
Regarding the sharpness of (4) and (5), we could not find anything in the literature. Thus, as our first theorem, we show that, just like in the one-parameter setting, both of these inequalities are indeed equivalences. Before doing so, let us briefly recall the notion of Carleson families, the John-Nirenberg lemma in the product setting, and a weak form of the Fefferman-Cordoba covering lemma for rectangles.
@@
Finally, it follows from the Fefferman-Cordoba covering lemma that, for any collection of rectangles $\mathcal{C}$, it is possible to extract a $\frac{1}{2}$-sparse sub-collection $\mathcal{C}^{\prime}$ such that
@@
See [9, 17] for a detailed exposition of these.
@@
The first case follows by plugging $f=h_{R}$ into the operator and using Lebesgue’s differentiation theorem. We now show that (5) is sharp. To this aim, normalize $g$ such that
@@
$\|S(f)\|_{L^{p}}^{2}=\|S(f)^{2}\|_{L^{\frac{p}{2}}}=\int S(f)^{2}\varphi,$
@@
$\|S(f)\|_{L^{p}}^{2}=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\chi_{R}\varphi=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\left\langle\varphi\right\rangle_{R}|R|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||R|.$
@@
$\|S(f)\|_{L^{p}}^{2}\lesssim\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||E_{R}|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\cup_{R\in\mathcal{C}_{k}}E_{R}}M(\varphi)\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\{M(g)>2^{k}\}}M(\varphi),$
@@
$\|S(f)\|_{L^{p}}^{2}\lesssim\int M(\varphi)M(g)^{2t}\leq\|M(\varphi)\|_{L^{(\frac{p}{2})^{\prime}}}\|M(g)\|_{L^{r}}^{\frac{2r}{p}}\lesssim\|M(g)\|_{L^{r}}^{\frac{2r}{p}},$
@@
(8) $\|S(F)\|_{L^{q}}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}}.$
@@
$\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{r}{q}}\lesssim\|S(F)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{r}{p}},$
@@ 3.1. The Operator $\pi^{3}_{g}$
and the above-mentioned mixed characterizations of product Hardy spaces [15, 16, 14, 17].
Unfortunately, we were unable to establish the sharpness of (9) in the full range of exponents. This is because the structure of the level sets of the involved mixed operators and their relationships is more complicated, and the picture is obscured. However, in Theorem 3.2, we show that when $1<r<\infty$, the lower bound follows directly from the one for the operator $\pi_{g}$ defined in (2).
Regarding the diagonal counterpart of (9), we show that the operator norm is comparable to a mixed-type norm of $g$, i.e.,
@@
which is simply the essential supremum of the $BMO_{d}(\mathbb{R})$-norm of the horizontal slices of $g$ (a similar notation is used for a mixture of $H^{r}$ and $L^{r}$ norms in (12)). See [15, 14], with the above quantity replaced by the $L^{\infty}(\mathbb{R}^{2})$-norm. We will establish this using the atomic decomposition theorem from [20]. To this aim, let us recall some simple but useful notions.
@@
Also, when $f$ is finite linear combination of rectangular Haar function, the this sum is finite. See [20] for the proof. In order to use this atomic decomposition we need to prove a simple lemma which is very useful when working with “local operators”.
@@
Next, we prove the upper bound for $\pi^{3}_{g}$, and to this aim again normalize $g$, such that
@@
To this aim, for a fixed $J$ consider the operator $\pi_{g^{\prime}}$ defined in (13) and note that since
@@ 3.2. The Operator $\pi^{4}_{g}$
and therefore, from boundedness of $\pi^{3}_{f^{\prime}}$ and duality it follows that
@@
$\vec{f}(x,y)_{J}:=f_{J}(x)|J|^{-\frac{1}{2}}\chi_{J}(x),\quad J\in\mathcal{D},\quad\vec{f^{\prime}}(x,y)_{I}:=f^{\prime}_{I}(y)|I|^{-\frac{1}{2}}\chi_{I}(x),\quad I\in\mathcal{D}.$
@@
and note that since the first $n$ generations of dyadic rectangles cover the unit square $n^{2}$ times we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq n$. Indeed since $g^{\prime}$ has a tensorial structure we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}=n$. Also, note that changing the signs of coefficients of $g^{\prime}$ does not change the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm of the function. Next, take an $n\times n$ matrix $\{H_{ij}:0\leq i,j<n\}$, with $\pm 1$ entries and with small norm. A perfect example is a Hadamard matrix, which is simply an orthogonal matrix with $\pm 1$ entries and therefore its operator norm on $\mathbb{R}^{n}$ equipped with the Euclidean norm is exactly $\sqrt{n}$. Such a matrix exists when $n$ is a power of 2. Now we modify $g^{\prime}$, obtain a new function $g$ defined by
@@
Interestingly, this example shows that the rectangular Haar basis is an unconditional basis for the space of functions equipped with the $\|.\|_{p}$-norm. So, one may wonder whether the remedy is to consider a stronger norm defined as
@@
However, a modification of the above example shows that, again, this quantity can be much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. To see this, simply let $H$ be the identity matrix, implying that $g_{I\times J}=1$ on the first $n$ generations of dyadic squares in $0,1]^{2}$. Since they cover the unit square $n$ times, we must have $\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq\sqrt{n}$. But now, at each point, the operator norm of the matrix $G(x,y)$ is $1$, and thus $\|g\|_{p}^{\prime}\lesssim 1$, again a quantity much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. Unfortunately, we were unable to combine this new argument with the previous ones invoking the atomic decomposition, and could not obtain an improved product BMO-type norm for $\pi^{4}_{g}$.
@@
To see this, let $\tilde{f}(x,y)=f(y,x)$, and then we have
@@
which combined with duality proves the above claim.
We conclude this note by mentioning a standard application of the lower bounds in weak factorization theorems for Hardy spaces. As an example, the bound
@@
and thus it follows directly from the general atomic decomposition in [8] that, for any function $g^{\prime}\in L^{r^{\prime}}(\mathbb{R}^{2})$, there exists a sequence of positive numbers $\{\lambda_{i}\}_{i\geq 0}$, and pairs of functions $\{f_{i},f^{\prime}_{i}\}_{i\geq 0}$ in the unit balls of $L^{p}(\mathbb{R}^{2})$ and $L^{q^{\prime}}(\mathbb{R}^{2})$, such that
@@
See [13] for similar applications of this result. On the other hand, from the above example,
$\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\not\simeq\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}|\left\langle\pi^{3}_{f^{\prime}}(f),g\right\rangle|,$
and thus there exists a function $g^{\prime}\in H_{d}^{1}(\mathbb{R}\otimes\mathbb{R})$ for which any factorization of the form
@@
would fail to hold. In particular, the Haar coefficients of $g^{\prime}$ cannot be of the form
$g^{\prime}_{I\times J}=\left\langle f_{J}\otimes f^{\prime}_{I}\right\rangle_{I\times J},\quad I,J\in\mathcal{D},\quad\sum_{I,J\in\mathcal{D}}\|f_{J}\otimes f^{\prime}_{I}\|_{L^{2}(\mathbb{R}^{2})}^{2}=1.$

### v2

@@ 1. Introduction
This work is a sequel to our previous two papers on the boundedness properties of paraproducts [24, 25]. Here, we are concerned with sharpness of the existing bounds for the remaining dyadic bi-parameter paraproducts not studied in [24]. For various applications and properties of paraproducts, we refer the reader to [3, 12, 13, 2, 19, 20, 21, 18].
@@
$\pi^{\varepsilon}(f,g)(x):=\sum_{I\in\mathcal{D}}\left\langle f,h_{I}^{\varepsilon_{1}}\right\rangle\left\langle g,h_{I}^{\varepsilon_{2}}\right\rangle h_{I}^{\varepsilon_{1}+\varepsilon_{2}}(x),\quad\varepsilon=(\varepsilon_{1},\varepsilon_{2})\in\{0,1\}^{2}\backslash\{(0,0)\},\quad x\in\mathbb{R}.$
In the above, $I$ stands for a dyadic interval on the line and $\mathcal{D}$ for the collection of all such intervals. Also, $h_{I}$ is the $L^{2}$-normalized Haar wavelet associated to $I$ and $h_{I}^{\epsilon}$ is given by
$h_{I}^{\epsilon}(x):=\begin{cases}h_{I}(x)&\epsilon=1\\ \bar{\chi}_{I}(x)&\epsilon=0,\end{cases}\quad x\in\mathbb{R},$
where $\bar{\chi}_{I}:=\frac{\chi_{I}}{|I|}$. Finally, the sum $\varepsilon_{1}+\varepsilon_{2}$ is understood in mode 2. It is then easy to see that for sufficiently nice functions we have
@@
Fairly well-known arguments show that, for $0<p<\infty$, the bilinear form $\pi^{(0,1)}$ is bounded from $H^{p}_{d}(\mathbb{R})\times BMO_{d}(\mathbb{R})$ and $L^{\infty}(\mathbb{R})\times H^{p}_{d}(\mathbb{R})$ to $H^{p}_{d}(\mathbb{R})$ [2, 12, 19, 20, 13, 21, 18]. Here, for $0<p<\infty$, $H^{p}_{d}(\mathbb{R})$ denotes the dyadic Hardy space on the line, (quasi-)normed by
@@
It is enough to recall the square function characterization of dyadic Hardy spaces, stating that
@@
As expected, the $H_{d}^{p}(\mathbb{R})$-to-$H_{d}^{q}(\mathbb{R})$ norm of $\pi^{\prime\prime}_{g}$ is also comparable to $\|g\|_{H_{d}^{r}(\mathbb{R})}$ in the full range of exponents, when $\frac{1}{q}=\frac{1}{p}+\frac{1}{r}$. In the reflexive range, this result was established in [22] (Theorem 12.2, p. 128), and in Theorem 2.1 we provide a proof of this in the product setting, to which we now turn.
@@ 2. Unmixed Paraproducts
where the supremum is taken over all open sets of finite positive measure. Since it will be clear from the context, we use the same notation $M$ for the dyadic maximal operator on the line and the strong operator defined above. We also refer the reader to [24] for the reason behind the above equivalences, and to [4, 7, 26, 11, 8] for more on Hardy spaces and BMO in the product setting. See also [6, 23]. The adjoint of this operator, which is unmixed too, is given by
@@
$\|f\|_{H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})}\simeq\|S(f)\|_{L^{p}(\mathbb{R}^{2})},\quad 0<p<\infty,\quad S(f):=\Big(\sum_{R}|f_{R}|^{2}\frac{\chi_{R}}{|R|}\Big)^{\frac{1}{2}},$
[7, 26]. Now, the above bounds are simple consequences of the crucial pointwise inequality
@@
the square function characterization of product Hardy spaces, and Hölder’s inequality [19, 20, 21, 18].
Regarding the sharpness of (4) and (5), we could not find anything in the literature. Thus, as our first task, we show that, just like in the one-parameter setting, both of these inequalities are indeed equivalences. Before doing so, let us briefly recall the notion of Carleson families, the John-Nirenberg lemma in the product setting, and a weak form of the Fefferman-Córdoba covering lemma for rectangles.
@@
Here, the case $0<p\leq 1$, follows simply from Hölder’s inequality and we refer the reader to [21] for the proof of the general case. Finally, it follows from the Fefferman-Córdoba covering lemma that, for any collection of rectangles $\mathcal{C}$, it is possible to extract a $\frac{1}{2}$-sparse sub-collection $\mathcal{C}^{\prime}$ such that
@@
See [10, 11, 21] for a detailed exposition of these.
@@
The first case follows by plugging $f=h_{R}$ into the operator and using Lebesgue’s differentiation theorem. We now show that (5) is sharp. To this end, normalize $g$ such that
@@
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{2}=\|S(f)^{2}\|_{L^{\frac{p}{2}}(\mathbb{R}^{2})}=\int S(f)^{2}\varphi,$
@@
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{2}=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\chi_{R}\varphi=\int\sum_{R\in\mathcal{C}}2^{2t\lambda(R)}\left\langle\varphi\right\rangle_{R}|R|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||R|.$
@@
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{2}\lesssim\sum_{k\in\mathbb{Z}}2^{2tk}\sum_{R\in\mathcal{C}_{k}}|\left\langle\varphi\right\rangle_{R}||E_{R}|\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\cup_{R\in\mathcal{C}_{k}}E_{R}}M(\varphi)\leq\sum_{k\in\mathbb{Z}}2^{2tk}\int_{\{M(g)>2^{k}\}}M(\varphi),$
@@
$\|S(f)\|_{L^{p}(\mathbb{R}^{2})}^{2}\lesssim\int M(\varphi)M(g)^{2t}\leq\|M(\varphi)\|_{L^{(\frac{p}{2})^{\prime}}(\mathbb{R}^{2})}\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{2r}{p}}\lesssim\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{2r}{p}},$
@@
(8) $\|S(F)\|_{L^{q}(\mathbb{R}^{2})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}}.$
@@
$\|M(g)\|_{L^{r}(\mathbb{R}^{2})}^{\frac{r}{q}}\lesssim\|S(F)\|_{L^{q}(\mathbb{R}^{2})}\lesssim\|g\|_{H_{d}^{r}(\mathbb{R}\otimes\mathbb{R})}^{\frac{r}{p}},$
@@ 3.1. The Operator $\pi^{3}_{g}$
and the above-mentioned mixed characterizations of product Hardy spaces [19, 20, 18, 21]. Regarding this it is natural to ask:
Question 3.1.
Is it true that (9) is an equivalence in the full range of exponents?
Theorem 3.3 answers this question in the positive in the reflexive range and provides a lower bound in the general case. However, the more intricate structure of the level sets of the mixed operators involved, and their interrelations, leaves the case $0<r\leq 1$ open. Nevertheless, the same theorem shows that for all $0<p<\infty$, the operator norm of $\pi^{3}_{g}$ on $H_{d}^{p}(\mathbb{R}\otimes\mathbb{R})$ is comparable to a mixed-type norm of $g$, i.e.
@@
which is simply the essential supremum of the $BMO_{d}(\mathbb{R})$-norm of the horizontal slices of $g$ (a similar notation is used for a mixture of $H^{r}$ and $L^{r}$ norms in (12)). See [19, 18], with the above quantity replaced by the $L^{\infty}(\mathbb{R}^{2})$-norm. We will establish this using the atomic decomposition theorem from [24]. To this end, let us recall some simple but useful notions.
@@
Also, when $f$ is finite linear combination of rectangular Haar functions this sum is finite. See [24] for the proof. In order to use this atomic decomposition we need to prove a simple lemma which is very useful when working with “local operators”.
@@
Next, we prove the upper bound for $\pi^{3}_{g}$, and to this end again normalize $g$, such that
@@
To this end, for a fixed $J$ consider the operator $\pi_{g^{\prime}}$ defined in (13) and note that since
@@ 3.2. The Operator $\pi^{4}_{g}$
and therefore, from the boundedness of $\pi^{3}_{f^{\prime}}$ and duality it follows that
@@
$\vec{f}(x,y)_{J}:=f_{J}(x)|J|^{-\frac{1}{2}}\chi_{J}(y),\quad J\in\mathcal{D},\quad\vec{f^{\prime}}(x,y)_{I}:=f^{\prime}_{I}(y)|I|^{-\frac{1}{2}}\chi_{I}(x),\quad I\in\mathcal{D}.$
@@
and note that the dyadic rectangles coming from the product of the first $n$ generations of dyadic intervals in $[0,1]$, cover the unit square $n^{2}$ times. Therefore, we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq n$. Indeed since $g^{\prime}$ has a tensorial structure we have $\|g^{\prime}\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}=n$. Also, note that changing the signs of coefficients of $g^{\prime}$ does not change the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm of the function. Next, take an $n\times n$ matrix $\{H_{ij}:0\leq i,j<n\}$, with $\pm 1$ entries and with small norm. A perfect example is a Hadamard matrix, which is simply an orthogonal matrix with $\pm 1$ entries and therefore its operator norm on $\mathbb{R}^{n}$ equipped with the Euclidean norm is exactly $\sqrt{n}$. Such a matrix exists when $n$ is a power of 2. Now we modify $g^{\prime}$, obtain a new function $g$ defined by
@@
Interestingly, this example shows that the rectangular Haar basis is not an unconditional basis for the space of functions equipped with the $\|.\|_{p}$-norm. So, one may wonder whether the remedy is to consider a stronger norm defined as
@@
However, a modification of the above example shows that, again, this quantity can be much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. To see this, simply let $H$ be the identity matrix, implying that $g_{I\times J}=|I\times J|^{\frac{1}{2}}$ on the first $n$ generations of dyadic squares in $0,1]^{2}$. Since they cover the unit square $n$ times, we must have $\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\geq\sqrt{n}$. But now, at each point, the operator norm of the matrix $G(x,y)$ is $1$, and thus $\|g\|_{p}^{\prime}\lesssim 1$, again a quantity much smaller than the $\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})$-norm. Unfortunately, we were unable to combine this new argument with the previous ones invoking the atomic decomposition, and could not obtain an improved product BMO-type bound for $\pi^{4}_{g}$. So let us pose the following question:
Question 3.5.
Is there a reasonable characterization of the operator norm of $\pi^{4}_{g}$? In particular, is there a way to combine the above matrix point of view with the atomic decomposition techniques to provide an improved bound for this operator?
@@
To see this, let $\tilde{f}(x,y)=f(y,x)$, and then we note that
@@
which implies $\|g\|_{r}=\|g\|_{r^{\prime}}$, for any $1<r<\infty$. Therefore, if we choose $1<r<\min\{p,p^{\prime},q,q^{\prime}\}$, we must have $\|g\|_{p}\simeq\|g\|_{q}\simeq\|g\|_{r}$, proving our claim.
We conclude this note by mentioning a standard application of the lower bounds in the weak factorization theorems for Hardy spaces. Here, a basic tool is a general atomic decomposition stating that:
Theorem.
Let $V\subset X$ be a bounded subset of a Banach space $X$, symmetric around the origin, and such that for every $l\in X^{*}$,
$\|l\|_{X^{*}}\simeq\sup_{x\in V}|l(x)|.$
Then, for every $x\in X$, there exists a sequence of positive real numbers $\{\lambda_{i}\}_{i\geq 1}$ such that
$x=\sum_{i\geq 1}\lambda_{i}v_{i},\quad\|x\|_{X}\simeq\sum_{i\geq 1}\lambda_{i},\quad v_{i}\in V,\quad i\geq 1.$
In the above, the series converges in the norm topology of $X$.
See [9, 17] for the proof and applications of this result. Now, as an example, the bound
@@
and thus it follows directly from the above theorem that, for any function $g^{\prime}\in L^{r^{\prime}}(\mathbb{R}^{2})$, there exists a sequence of positive numbers $\{\lambda_{i}\}_{i\geq 0}$, and pairs of functions $\{f_{i},f^{\prime}_{i}\}_{i\geq 0}$ in the unit balls of $L^{p}(\mathbb{R}^{2})$ and $L^{q^{\prime}}(\mathbb{R}^{2})$, such that
@@
On the other hand, the converse statement to the above theorem holds as well, and as our earlier example shows
$\|g\|_{\text{BMO}_{d}(\mathbb{R}\otimes\mathbb{R})}\not\simeq\sup_{\begin{subarray}{c}\|f\|_{L^{p}(\mathbb{R}^{2})}=1\\ \|f^{\prime}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}=1\end{subarray}}|\left\langle\pi^{3}_{f^{\prime}}(f),g\right\rangle|.$
Thus, there exists a function $g^{\prime}\in H_{d}^{1}(\mathbb{R}\otimes\mathbb{R})$ for which any factorization of the form
@@
would fail to hold. In particular, $g^{\prime}$ is not of the form $\pi^{3}_{f^{\prime}}(f)$, for any pair of functions in the unit ball of $L^{2}(\mathbb{R}^{2})$. Equivalently, for any two families of functions on line $\{f_{J}\}_{J\in\mathcal{D}}$ and $\{f^{\prime}_{I}\}_{I\in\mathcal{D}}$ satisfying
$\sum_{J\in\mathcal{D}}\|f_{J}\|_{L^{2}(\mathbb{R}^{2})}^{2}=1=\sum_{I\in\mathcal{D}}\|f^{\prime}_{I}\|_{L^{2}(\mathbb{R}^{2})}^{2},$
the Haar coefficients of $g^{\prime}$ cannot be of the form
$g^{\prime}_{I\times J}=\left\langle f_{J}\otimes f^{\prime}_{I}\right\rangle_{I\times J},\quad I,J\in\mathcal{D}.$
Acknowledgments
I would like to thank the referee for valuable comments on the manuscript.